# The Artificial Scientist: in-transit Machine Learning of Plasma Simulations

### v2

@@ Abstract
Large-scale simulations or scientific experiments produce petabytes of data per run. This poses massive challenges for I/O and storage when scientific analysis workflows are run manually offline. Unsupervised deep learning-based techniques to extract patterns and non-linear relations from these large amounts of data provide a way to build scientific understanding from raw data, reducing the need for manual pre-selection of analysis steps, but require exascale compute and memory to process the full dataset available. In this paper, we demonstrate a heterogeneous streaming workflow in which plasma simulation data is streamed directly to a Machine Learning (ML) application training a model on the simulation data in-transit, completely circumventing the capacity-constrained filesystem bottleneck. This workflow employs openPMD to provide a high level interface to describe scientific data and also uses ADIOS2, to transfer volumes of data that exceed the capabilities of the filesystem. We employ experience replay to avoid catastrophic forgetting in learning from this non-steady state process in a continual manner and adapt it to improve model convergence while learning in-transit. As a proof-of-concept we approach the ill-posed inverse problem of predicting particle dynamics from radiation in a particle-in-cell (PIConGPU) simulation of the Kelvin-Helmholtz instability (KHI). We detail hardware-software co-design challenges as we scale PIConGPU to full Frontier, the Top-1 system as of June 2024 Top500 list.
@@
HPC simulations capable of fully exploiting the compute power of supercomputers produce a wealth of high quality, complex data. Examples for such simulations include digital twins of the earth, whole device models of magnetic fusion reactors or simulations of advanced particle accelerators. The data rates and volumes of such simulations can be compared to instruments at large-scale research infrastructures such as tomography endstations at synchrotrons in photon science, high energy physics experiments, fusion research facilities or astro-particle observatories [1].
Exploiting such a huge volume of data optimally often focuses on fast data selection including data filtering, triggering, anomaly and outlier detection or intelligent data reduction. In-situ inference using pre-trained machine learning (ML) models has led to a broad inclusion of such techniques across scientific communities [2, 3, 4, 5].
@@
Here we want to highlight an open challenge, namely the situation in which the stream of data produced by a simulation or any other source such as a scientific instrument remains at such a scale and quality that data reduction still demands in-transit analysis. In such cases, even if the volume of data is reduced to the minimum extent suitable for analysis, storage of such data for further offline analysis is impossible due to the physical constraints of the filesystem at hand.
@@
The open challenge in our case is a particle-in-cell simulation of a Kelvin-Helmholtz instability (KHI) [8] as shown in Fig. 1 (more details in Section II). Instabilities are one of the most fundamental yet complex collective phenomena in plasma physics. Their growth rate, scale lengths and dynamics show a strong non-linear dependence on the plasma properties. Understanding and controlling instabilities is critical across all disciplines of plasma physics from magnetic and inertial fusion to space physics, plasma acceleration or astrophysics. Ultimately, reconstructing the complete phase space dynamics of a plasma from observations would provide us with the most detailed view of the growth and dynamics of instabilities and thus on the most central plasma properties at high temporal and spatial detail.
@@
Reconstructing the complete phase space dynamics of a plasma requires processing of a wealth of information. One of the most challenging use cases in this class of problems is the in-transit training of data-driven complex machine learning models at extreme data rates, where training the model requires most of the data produced by the source. The model being trained itself is of such complexity that it can in principle incorporate the complexity of the total volume of data produced by the source, while at no time a total view of the data is available.
@@
We note that a large number of advances in ML rely on iterative training [9], where the model repeatedly learns from the input data over multiple epochs, requiring permanent storage of the datasets. In the reference case studied in our work, this is not the case, as data is produced on demand and discarded after being used for training. We thus have chosen a continuous learning approach to address this challenge. Although methods for continuous learning exist [10], they have not had to consider the time constraints imposed by the speed of data generated from simulations [11]. Thus, adapting the continuous learning techniques to handle such constraints is a critical challenge in the field of ML for high rate streaming of scientific data.
@@
The other challenge is to effectively navigate the memory hierarchy of heterogeneous computing systems, as it exposes enormous differences in throughput [12, 13]. While internal throughput in GPUs can easily reach the TB/s range on a single node, breaking down the throughput of massively parallel filesystems to the single node throughput reveals maximum performance of several tens of MB/s only. Each node often houses more than one GPU, hence available data storage and bandwidth demand data reduction, often done by discarding highly valuable data in practice.
@@
We further note that the mismatch in throughput between memory hierarchies in modern HPC systems has been steadily growing, increasing the number of cases in which offline analysis is impossible and thus the urgency to find solutions for online analysis of data at high rates and volumes [12]. We argue that in case of such extreme rates online analysis of streaming data is the method of choice [14, 7], reducing the overall amount of data that needs to be stored while using computationally demanding algorithms to make optimum use of the data produced. However, this means that a complete view of all data produced is never available at any time of the analysis workflow and there is a need to preserve the extreme data throughput during the whole analytics pipeline.
@@
In our work we showcase a real-world example and present a general solution that implements all relevant parts of online in-situ training of a large-scale model on the Frontier exascale supercomputer at ORNL [15]. We argue that the building blocks of our solution are general enough to be adapted to almost any use case of the type discussed above with only minor adjustments. We furthermore emphasize that besides presenting a full workflow, the additional unique value of this work is to highlight the multi-parameter optimization problem of fitting the available compute and data transfer resources to achieve the best result in the shortest amount of time. We remark that our technique is especially interesting in the case of irreversible data loss [16], and thus extends well beyond the case of large-scale simulations as high rate data sources presented here.
@@
Scientifically: We uncover correlations between emitted radiation and particle dynamics within the simulation in an unsupervised manner to create an inverse map from the observed radiation to the particle dynamics and to identify the aspects of the particle dynamics that are relevant to the remotely observable radiation signatures of KHI.
@@ II The relativistic Kelvin-Helmholtz instability
The Kelvin-Helmholtz instability (KHI) is a well-known shear surface instability observed in plasmas [8]. It is driven by a self-amplifying cycle of small density or velocity fluctuations that lead to a growing magnetic field at the shear surface, which further amplifies the initial fluctuations as depicted in Fig. 1. This phenomenon occurs in astrophysical scenarios, including Saturn’s magnetopause [18], but also in fusion reactors and inertial confinement fusion (ICF) [19, 20]. While the theoretical understanding and simulation capabilities of KHI [21, 22] are well established, direct measurement of the predicted density changes remains a real challenge. Because of its well-understood dynamics, we chose KHI as a test case for our study to validate the results of our orchestrated workflow against existing knowledge.
@@
Here, we simulate the KHI using the plasma simulation code PIConGPU [24, 6] which utilizes the alpaka library [25] for performance portability and has been at the forefront of developing central technologies for GPU-accelerated plasma simulations. It was the first particle-in-cell (PIC) [26] code to run performant simulations on a variety of architectures in the world over many years [7].
@@ III Reconstructing the local phase space dynamics of the relativistic KHI from observable radiation
Fig. 2 depicts the interrelations between observational data, such as radiation, and simulation state, such as particles, as both inputs and outputs to data-driven models. Our objective of approximating the inverse mapping from observed radiation to particle dynamics can abstractly be described by the information flow depicted in Fig. 2(a). For the model to remain tractable a reduced representation of the complex particle dynamics is required, which is learned by an auto-encoding model Fig. 2(b). To observe the extent to which this representation captures relevant features of the particle dynamics, it is worthwhile to train a surrogate model of the, computationally expensive, forward process of radiation emission (Fig. 2(c)). These three tasks integrate into our ML model, described later, also in Fig. 7.
It was shown in 2013 [6] that the computation of high quality, high resolution observational data such as radiation spectra can only be performed in-situ, as storing the ground truth data, the phase space trajectory data of each particle in the simulation at each time step, is impossible due to the aforementioned throughput hierarchy. As an example, scaling to a moderate 25 % of the Frontier system, one would be faced with 1 PB of data for every time step. With total time steps per simulation of the order of a few thousand time steps and time step durations of the order of 0.1 to 1 s, we frequently observe data rates of 1 to 10 PB/s for particle trajectory data and theoretically require total volumes on disk of the order of 10 EB. Depleting storage in as little as . We thus end up with the conundrum to employ data-driven models for solving a large-scale ill-posed inverse problem.
@@ III-A A complex in-memory workflow to reconstruct local phase space dynamics from radiation
A particular challenge for the reconstruction of local phase space dynamics is that their radiation signatures are highly convoluted, composite quantities that combine multiple facets of the ground truth which to restore. This observation guides our approach at learning from a live PIC simulation, as the posed problem is now diagnosed as ill-posed. In detail, the observed radiation spectrum is both spectrally and angularly resolved and the observed intensity at each wavelength and direction varies between time steps according to the changes of the phase space dynamics. The observed intensity in each direction and at each wavelength is a weighted sum of all contributions from all particles.
Hence here are the steps to be taken:
Simulate a sufficiently large domain at high spatial and temporal resolution to capture both the local plasma dynamics as well as the collective contributions of different plasma regions to the total radiation spectrum.
@@
Calculate the radiation spectrum directly in-situ as the sum of the contributions from all radiating particles in the simulation at each time step, with high angular and spectral resolution and bandwidth.
Collect the particle phase space data, positions and momenta, and the complex amplitude and phase for each wavelength and direction of the radiation.
Prepare the collected data for an ML model by finding a suitable encodings for spectral and phase space data.
Stream that data to an ML model and asynchronously train the model.
Stop the simulation after observing the complete physical process under investigation.
@@ III-B The technical challenge: Exascale data streaming and continual learning
An important design goal for the implementation of the above data processing scheme is to write no simulation data to the filesystem during the entire workflow. File I/O can certainly be initiated when desired, but is independent of and not necessary for the workflow.
While in-situ approaches have been previously used to compute scientifically relevant analyses through plugins within the same application context [6], this is no longer possible for the envisioned ML analysis, as the vastly incompatible software stacks for the C++ simulation code PIConGPU and the PyTorch-based MLapp as well as their different scaling properties inhibit such an integration into a strongly-coupled monolithic application framework. Hence, simulation and the MLapp are modularized into two separate applications that communicate through a standardized data interface, a pattern that we refer to as loose coupling, since the standardized data interface allows to plug together a data processing pipeline more flexibly, i.e. loosely than before. When the implementation for such workflows avoids writing intermediate data to disk, we refer to them as in-memory. In-memory workflows include such setups where data is not moved, staying within one application’s memory (in-situ), and alternatively setups that move data from one into another application’s memory (in-transit) (Fig. 3(a)). The latter becomes necessary when transitioning from file-based workflows to in-memory workflows without sacrificing flexibility in the data exchange pattern, as well as in response to codes that scale differently.
The whole simulation, all data produced by the simulation and ingested by the model, and the whole model all reside in-memory and data is distributed and optionally reduced in-transit (Fig. 3(b)).
@@
Start a scalable, GPU-accelerated PIConGPU simulation of sufficient size and resolution, utilizing a large subset of nodes on Frontier.
Schedule PyTorch and N/RCCL alongside the PIConGPU simulation. (N/RCCL - NVIDIA and ROCm collective communication library)
@@
At each simulation time step
stream particle momenta and positions and spectral data to a large-scale ML model,
@@
For supporting an online workflow which stores no intermediate data such as phase space trajectory and radiation data to disk, the model must be continuously trained online from the series of time step data. This requires the model to be trained by subsequent snapshots of the evolution of plasma and radiation data. For this, continual learning is the methodology of choice as outlined in Section IV-C.
We strongly believe that a large class of inversion problems for complex, large-scale systems that produce data streams face similar challenges and that our solution presented here for the first time introduces all relevant building blocks and technologies to address these challenges.
@@ IV-A Particle-in-cell simulation and its setup
On Frontier with a total of AMD MI250X GPUs across nodes [28], a scaling of PIConGPU’s Figure of Merit (FOM), i. e. the weighted sum of the total number of particle updates per second (90 %) and the number of cell updates per second (10 %), is demonstrated as shown in the weak scaling results. In this particular configuration, the simulation included macroparticles in grid cells, compared to our base FOM run on Summit in 2019, where only macroparticles in cells could be used.
@@ IV-B Data streaming via openPMD
Considering that PIConGPU as well as the MLapp both rely significantly on very different software stacks, launching both parts of the setup in the same application context is not feasible. As detailed previously in III-B, we follow a loose-coupling rather than a tight-coupling approach [14], launching both applications separately. Output created by PIConGPU must be moved across the application boundary to the MLapp, without impacting the scalability of the setup through use of limited resources such as the parallel filesystem. The solution pursued in this paper is an in-transit workflow that keeps data in memory at all times and writes no intermediate data to disk.
@@
The in-transit data strategy is built upon the Sustainable Staging Transport (SST) [39, 40] data engine of ADIOS2 that connects one parallel data producer code to an arbitrary number of parallel data consumer codes, hereby opening connections only between processes that actually share data. For adapting to different system architectures, the SST engine implements different network transport technologies (data planes), including TCP (non-scalable fallback), libfabric, ucx and the MPI_Open_port() API of MPI. For this benchmark, we consider the low-level libfabric data plane on top of the system’s CXI provider [41] which allows, yet also requires more manual finetuning, as well the MPI data plane which leverages the MPI implementation’s performance tuning.
@@
The PIConGPU KHI simulation is run from half (4096 compute nodes) to full-scale (9126 compute nodes) of Frontier, producing of particle data per compute node and time step. Fig. 3(c) (bottom left) shows the processes running within each compute node and the data flow between them. At each scaling run (weak scaling), five time steps are sent from PIConGPU to the no-op consumer, which then measures the time needed for loading the data. The perceived parallel throughput is computed based on this measured time and the global data size, i. e. the previously mentioned per node multiplied with the number of nodes. While including communication overhead, this perceived parallel throughput has been shown in [40] to be a close approximation for the real throughput in this kind of setup.
The benchmarks in Fig. 6 demonstrate the feasibility of this I/O strategy for full-scale workflows. Depicted are boxplots of all single measurements. Most notably, we observe a maximum perceived parallel throughput of 20 - 30 TB/s which compares outstandingly against the 10 TB/s bandwidth of the parallel Orion filesystem, a scaling limit which we circumvent and exceed by not using the filesystem for intermediate data. Furthermore, these results can also compare well against the 35 TB/s aggregate write bandwidth of the SSDs installed locally in the compute nodes [42].
For the libfabric-based benchmarks, we initially used an implementation that enqueued all read operations to the network at once and waited for replies (labeled in Fig. 6). While these results signify the best achieved per-node throughput of 3.5 - 4.7 GB/s at 4096 compute nodes (a total of 14.1 - 18.8 TB/s), it turned out that this strategy did not scale to the full system, so we employed an alternative that enqueued operations in batches of 10 operations. While this turned out to scale to the full system, it came at a notable cost in performance, visible in a per-node throughput of 1.9 - 2.6 GB/s at 9126 compute nodes (a total of 16.5 - 23.0 TB/s). Conversely, the MPI data plane yields a per-node throughput from 2.6 - 3.7 GB/s at 4096 compute nodes (a total of 10.5 - 14.9 TB/s) to 2.4 - 3.3 GB/s at 9126 nodes (a total of 21.4 - 29.5 TB/s), the best achieved system-wide parallel throughput from this experiment.

### v3

@@ Abstract
Large-scale simulations or scientific experiments produce petabytes of data per run. This poses massive challenges for I/O and storage when scientific analysis workflows are run manually offline. Unsupervised deep learning-based techniques to extract patterns and non-linear relations from these large amounts of data provide a way to build scientific understanding from raw data, reducing the need for manual pre-selection of analysis steps, but require exascale compute and memory to process the full dataset available. In this paper, we demonstrate a heterogeneous streaming workflow in which plasma simulation data is streamed directly to a Machine Learning (ML) application training a model on the simulation data in-transit, completely circumventing the capacity-constrained filesystem bottleneck. This workflow employs openPMD to provide a high level interface to describe scientific data and also uses ADIOS2, to transfer volumes of data that exceed the capabilities of the filesystem. We employ experience replay to avoid catastrophic forgetting in learning from this non-steady state process in a continual manner and adapt it to improve model convergence while learning in-transit. As a proof-of-concept, we approach the ill-posed inverse problem of predicting particle dynamics from radiation in a particle-in-cell (PIConGPU) simulation of the Kelvin-Helmholtz instability (KHI). We detail hardware-software co-design challenges as we scale PIConGPU to full Frontier, the Top-1 system as of June 2024 Top500 list.
@@
HPC simulations capable of fully exploiting the compute power of supercomputers produce a wealth of high quality, complex data. Examples for such simulations include digital twins of the earth, whole device models of magnetic fusion reactors or simulations of advanced particle accelerators. The data rates and volumes of such simulations can be compared to instruments at large-scale research infrastructures such as tomography endstations at synchrotrons in photon science, high energy physics experiments, fusion research facilities or astro-particle observatories [1]. Exploiting such a huge volume of data optimally often focuses on fast data selection including data filtering, triggering, anomaly and outlier detection or intelligent data reduction. In-situ inference using pre-trained machine learning (ML) models has led to a broad inclusion of such techniques across scientific communities [2, 3, 4, 5].
@@
Here we want to highlight an open challenge, namely the situation where a stream of data produced by a source – like a large-scale simulation or a high-resolution and high-repetition-rate experiment – is of such a scale and quality that, even after simple data reduction, storage for further offline analysis is impossible due to the physical bandwidth or capacity constraints of the filesystem at hand. In such cases data reduction already demands complex in-transit analysis.
@@
We illustrate this challenge with a particle-in-cell simulation of a Kelvin-Helmholtz instability (KHI) [8] as shown in Fig. 1 (more details in Section II). Instabilities are one of the most fundamental, yet complex collective phenomena in plasma physics. Understanding and controlling them is critical across all disciplines of plasma physics from magnetic and inertial fusion to space physics, plasma acceleration or astrophysics. Instability growth rate, length scales and dynamics show a strong non-linear dependence on the plasma properties. Reconstructing the complete phase space dynamics of a plasma from observations would provide the most detailed view of the growth and dynamics of instabilities, and thus of the most central plasma properties at high temporal and spatial detail.
@@
Reconstructing the complete phase space dynamics of a plasma requires processing of a wealth of information. One of the most challenging use cases in this class of problems is the in-transit training of data-driven complex ML models at extreme data rates, where training the model requires most of the data produced by the source. The model being trained itself must be of sufficient complexity that it can, in principle, incorporate the complexity of the total volume of data produced by the source, while at no time a total view of the data is available.
@@
We note that a large number of advances in ML rely on iterative training [9], where the model repeatedly learns from the input data over multiple epochs, requiring permanent storage of the datasets. In the reference case studied in our work, this is not the case, as data is produced on demand and discarded after being used for training. Due to the size of data used for training and memory constraints, no total view of the training data is available at any given time. We thus have chosen a continuous learning approach to address this challenge. Although methods for continuous learning exist [10], they have not had to consider the time constraints imposed by the speed of data generated from simulations [11]. Thus, adapting the continuous learning techniques to handle such constraints is a critical challenge in the field of ML for high rate streaming of scientific data.
@@
A challenge to scaling is to effectively navigate the memory hierarchy of heterogeneous computing systems, as it exposes enormous differences in throughput [12, 13]. While internal throughput in GPUs can easily reach the TB/s range on a single node, breaking down the throughput of massively parallel filesystems to the single node throughput reveals maximum performance of several tens of MB/s only. Each node often houses more than one GPU, hence available data storage and bandwidth demand data reduction, often done by discarding highly valuable data in practice.
@@
We further note that mismatch in throughputs between memory hierarchies in modern HPC systems has been steadily growing, increasing the number of cases in which offline analysis is impossible and thus the urgency to find solutions for online analysis of data at high rates and volumes [12, 14]. We argue that in case of such extreme rates online analysis of streaming data is the method of choice [15, 7], reducing the overall amount of data that needs to be stored while using computationally demanding algorithms to make optimum use of the data produced. However, this means that a complete view of all data produced is never available at any time of the analysis workflow and there is a need to preserve the extreme data throughput during the whole analytics pipeline.
@@
In our work we showcase a real-world example and present a general solution that implements all relevant parts of online in-situ training of a large-scale model on the Frontier exascale supercomputer at ORNL [16]. We argue that the building blocks of our solution are general enough to be adapted to almost any use case of the type discussed above with only minor adjustments. We furthermore emphasize that besides presenting a full workflow, the additional unique value of this work is to highlight the multi-parameter optimization problem of fitting the available compute and data transfer resources to achieve the best result in the shortest amount of time. We remark that our technique is especially interesting in the case of irreversible data loss [17]. It thus extends well beyond the extreme case of large-scale simulations presented here. Examples include other high data rate sources, e. g. plasma or particle experiments with high-resolution and high-repetition-rate detectors, which are more likely connected to smaller edge clusters.
@@
Scientifically: We uncover correlations between emitted radiation and particle dynamics within the simulation in an unsupervised manner to create an inverse map from the observed radiation to the particle dynamics, and to identify the aspects of the particle dynamics that are relevant to the remotely observable radiation signatures of KHI.
@@ II The relativistic Kelvin-Helmholtz instability
The Kelvin-Helmholtz instability (KHI) is a well-known shear surface instability observed in plasmas [8]. It is driven by a self-amplifying cycle of small density or velocity fluctuations that lead to a growing magnetic field at the shear surface, which further amplifies the initial fluctuations (as depicted in Fig. 1). This phenomenon occurs in astrophysical scenarios, including Saturn’s magnetopause [19], but also in fusion reactors and inertial confinement fusion (ICF) [20, 21]. While the theoretical understanding and simulation capabilities of KHI [22, 23] are well established, a direct measurement of the predicted density changes remains a real challenge. Because of its well-understood dynamics, we chose KHI as a test case for our study to validate the results of our orchestrated workflow against existing knowledge.
@@
Here, we simulate the KHI using the plasma simulation code PIConGPU [25, 6] which has been at the forefront of developing central technologies for GPU-accelerated plasma simulations, utilizing the alpaka library [26] for performance portability. It was the first particle-in-cell (PIC) [27] code to run performant simulations on a variety of architectures in the world over many years [7].
@@ III Reconstructing the local phase space dynamics of the relativistic KHI from observable radiation
Fig. 2 depicts the interrelations between observational data (such as radiation) and simulation state (such as particles) as both inputs and outputs to data-driven models. Our objective of approximating the inverse mapping from observed radiation to particle dynamics can abstractly be described by the information flow depicted in Fig. 2(a). For the model to remain tractable a reduced representation of the complex particle dynamics is required, which is learned by an auto-encoding model Fig. 2(b). To observe the extent to which this representation captures relevant features of the particle dynamics, it is worthwhile to train a surrogate model of the computationally expensive, forward process of radiation emission (Fig. 2(c)). These three tasks integrate into our ML model, described later, also in Fig. 7.
It was shown in 2013 [6] that the computation of high quality, high resolution observational data such as radiation spectra can only be performed in-situ, as storing the ground truth data, the phase space trajectory data of each particle in the simulation at each time step, is impossible due to the aforementioned throughput hierarchy. As an example, scaling to even a moderate 25 % of the Frontier system, one would be faced with 1 PB of data for every time step. With total time steps per simulation of the order of a few thousand time steps and time step durations of the order of 0.1 to , we frequently observe data rates of 1 to 10 PB/s for particle trajectory data and theoretically require total volumes on disk of the order of 10 EB. This means the storage could get depleted for a simulation that is as short as in duration. We thus end up with the conundrum to employ data-driven models for solving a large-scale ill-posed inverse problem without sufficient memory.
@@ III-A A complex in-memory workflow to reconstruct local phase space dynamics from radiation
A particular challenge for the reconstruction of local phase space dynamics is that their radiation signatures are highly convoluted, composite quantities that combine multiple facets of the ground truth to be restored. This observation guides our approach at learning from a live PIC simulation, as the posed problem is now diagnosed as ill-posed [28]. In detail, the observed radiation spectrum is both spectrally and angularly resolved and the observed intensity at each wavelength and direction varies between time steps according to the changes of the phase space dynamics. The observed intensity in each direction and at each wavelength is a weighted sum of all contributions from all particles.
Hence, here are the steps to be taken:
Simulate a sufficiently large domain at a high spatial and temporal resolution to capture both the local plasma dynamics, as well as the collective contributions of different plasma regions to the total radiation spectrum.
@@
Calculate the radiation spectrum in-situ as the sum of the contributions from all radiating particles in the simulation at each time step, with high angular and spectral resolution and high bandwidth.
Collect the particle phase space data (namely the positions and momenta) and the complex amplitude and phase for each wavelength and direction of the radiation.
Prepare the collected data for an ML model by finding suitable encodings for spectral and phase space data.
Stream the encoded data to an ML model and asynchronously train the model.
Stop the simulation after observing and covering a sufficient amount of relevant stages of plasma instability simulation.
@@ III-B The technical challenge: Exascale data streaming and continual learning
An important design goal for the implementation of the above data processing scheme is to write no simulation data to the filesystem during the entire workflow. File I/O can certainly be initiated when desired, but is not necessary for the workflow.
In-situ approaches have been previously used to compute scientifically relevant analyses through plugins within the same application context [6]. This is no longer possible for the envisioned ML analysis, as the vastly incompatible software stacks for the C++ simulation code PIConGPU and the PyTorch-based MLapp, as well as their different scaling properties inhibit such an integration into a strongly-coupled monolithic application framework. Hence, the simulation and the MLapp are modularized into two separate applications that communicate through a standardized data interface – a pattern that we refer to as loose coupling, since the standardized data interface allows to plug together a data processing pipeline more flexibly, i.e. loosely than before. When the implementation for such workflows avoids writing intermediate data to disk, we refer to them as in-memory. In-memory workflows include such setups where data is not moved, staying within one application’s memory (in-situ), and alternatively setups that move data from one into another application’s memory (in-transit) (Fig. 3(a)). The latter becomes necessary when transitioning from file-based workflows to in-memory workflows without sacrificing flexibility in the data exchange pattern, as well as in response to codes that scale differently.
The whole simulation, all data produced by the simulation and ingested by the model, and the whole model, all reside in-memory and data is distributed and optionally reduced in-transit (Fig. 3(b)).
@@
Start a scalable, GPU-accelerated PIConGPU simulation of sufficient size and resolution, utilizing a large subset of nodes on Frontier [29].
Schedule PyTorch and N/RCCL (i. e. NVIDIA and ROCm collective communication library) alongside the PIConGPU simulation.
@@
At each simulation time step:
Stream particle momenta, particle positions and the spectral data to a large-scale ML model,
@@
For supporting an online workflow which stores no intermediate data such as phase space trajectory and radiation data to disk, the model must be continuously trained online from the series of time step. This requires the model to be trained by subsequent snapshots of the evolution of plasma and radiation data. For this, continual learning is the methodology of choice as outlined in Section IV-C.
We strongly believe that a large class of inversion problems for complex, large-scale systems that produce data streams face similar challenges, and that our solution presented here for the first time introduces all the relevant building blocks and technologies to address these challenges.
@@ IV-A Particle-in-cell simulation and its setup
On Frontier with a total of AMD MI250X GPUs across nodes [29], a scaling of PIConGPU’s Figure of Merit (FOM), i. e. the weighted sum of the total number of particle updates per second (90 %) and the number of cell updates per second (10 %), is demonstrated as shown in the weak scaling results. For the results presented in this paper, the simulation involved macroparticles in grid cells, compared to our base FOM run on Summit in 2019, where only macroparticles in cells could be used [31].
@@ IV-B Data streaming via openPMD
Considering that both PIConGPU as well as the MLapp rely significantly on very different software stacks, launching both parts of the setup in the same application context is not feasible. As detailed previously in III-B, we follow a loose-coupling rather than a tight-coupling approach [15], launching both applications separately. Output created by PIConGPU must be moved across the application boundary to the MLapp, without impacting the scalability of the setup through use of limited resources such as the parallel filesystem. The solution pursued in this paper is an in-transit workflow that keeps data in memory at all times and writes no intermediate data to disk.
@@
The in-transit data strategy is built upon the Sustainable Staging Transport (SST) [42, 43] data engine of ADIOS2 that connects one parallel data producer code to an arbitrary number of parallel data consumer codes, thereby opening connections only between processes that actually share data. For adapting to different system architectures, the SST engine implements different network transport technologies (data planes), including TCP (non-scalable fallback), libfabric, ucx and the MPI_Open_port() API of MPI. For this benchmark, we consider the low-level libfabric data plane on top of the system’s CXI provider [44] which allows, yet also requires more manual finetuning, as well the MPI data plane which leverages the MPI implementation’s performance tuning.
@@
The PIConGPU KHI simulation is run from half (4096 compute nodes) to full-scale (9126 compute nodes) of Frontier, producing of particle data per compute node and time step. Fig. 3(c) (bottom left) shows the processes running within each compute node and the data flow between them. At each scaling run (weak scaling), five time steps are sent from PIConGPU to the no-op consumer, which then measures the time needed for loading the data. The parallel throughput is calculated based on this measured time and the global data size, i. e. the previously mentioned per node multiplied with the number of nodes. Although including communication overhead, this throughput value has been shown in [43] to be a close approximation for the real throughput in this kind of setup.
The benchmarks in Fig. 6 demonstrate the feasibility of this I/O strategy for full-scale workflows. Depicted are boxplots of all single measurements. Most notably, we observe a maximum parallel throughput of 20 - 30 TB/s which compares outstandingly against the 10 TB/s bandwidth of the parallel Orion filesystem, a scaling limit which we circumvent and exceed by not using the filesystem for intermediate data. Furthermore, these results can also compare well against the 35 TB/s aggregate write bandwidth of the SSDs installed locally in the compute nodes [45].
For the libfabric-based benchmarks, we initially used an implementation that enqueued all read operations to the network at once and waited for replies (labeled in Fig. 6). While these results signify the best achieved per-node throughput of 3.5 - 4.7 GB/s at 4096 compute nodes (a total of 14.1 - 18.8 TB/s), it turned out that this strategy did not scale to the full system. Thus, we employed an alternative that enqueued operations in batches of 10 operations. While this turned out to scale to the full system, it came at a notable cost in performance, visible in a per-node throughput of 1.9 - 2.6 GB/s at 9126 compute nodes (a total of 16.5 - 23.0 TB/s). Conversely, the MPI data plane yields a per-node throughput from 2.6 - 3.7 GB/s at 4096 compute nodes (a total of 10.5 - 14.9 TB/s) to 2.4 - 3.3 GB/s at 9126 nodes (a total of 21.4 - 29.5 TB/s), the best achieved system-wide parallel throughput from this experiment.